# Test-Time Training Scaling Laws for Chemical Exploration in Drug Design

### v2

@@ 1 Introduction
Thorough exploration of chemical space is critical to maximize the potential of generative AI in drug design. Identifying structurally diverse molecules with similar bioactivity increases the likelihood of discovering backup series with distinct chemical and biological profiles, mitigating risks in drug development [4]. Moreover, exploring novel chemical regions can circumvent patented compounds [5]. However, many generative models, including CLMs, suffer from mode collapse when optimizing for predicted properties, particularly those reliant on machine learning models with limited applicability domains due to epistemic uncertainty [6, 7].
@@ 3.1 Problem definition
However, chemical space is vast, property predictors are inherently inaccurate with high uncertainty, and Pareto optimal solutions have different property profiles. Hence, multiple molecules exist across chemical space that are considered hits if the score is above a given threshold (or sometimes more simply the top- molecules). We propose that the problem of generative molecular design should extend to finding all possible hits across chemical space, and postulate that this better aligns with practical requirements in drug discovery.
We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, even structurally unrelated molecular targets in chemical space (Figure 1, right). Firstly, previously published benchmarks typically don’t capture this with objectives focused on a single narrow target, or a single broad target (Figure 1, left). These do not adequately test the ability of an algorithm to explore chemical space if it is already receiving high rewards. Secondly, there are post-hoc metrics that measure the diversity of solutions [29, 30], however, this only measures the diversity inherently embedded in the objective. Therefore, designing an objective to contain multiple distinct targets provides a greater opportunity to observe inherent exploration capability. Lastly, previous benchmarks predominantly measure performance on the maximum achieved, whereas, we propose to measure performance on maximum for each molecular target representing a maxima in the reward landscape. This focuses evaluation on the algorithm’s ability to explore all high-rewarding regions.
@@ 3.2 Reinforcement learning with chemical language models
CLMs utilize a language representation, commonly SMILES [8], to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models. For example, SMILES represent the molecular structure of paracetamol with the string “CC(=O)Nc1ccc(O)cc1". These models operate autoregressively, predicting the next token in a sequence based on the context provided by previous tokens. The process can be formalized as maximizing the conditional probability of each token given the preceding sequence , where denotes the tokens up to . The model is trained using the cross-entropy loss function which is equivalent to the negative log-likelihood Equation 3. Where is the length of the -th sequence, and are the model parameters.
@@
We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the prior policy with a coefficient and by adding an exponent , as proposed by Thomas et al. [19]. These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape. In addition to reshaping the reward, we implement experience replay to augment on-policy data with off-policy data at each iteration. A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward. This additionally improves learning efficiency [37, 19].
@@ 3.3 Cooperative reinforcement learning
In the independent case, each agent still aims to maximize its cumulative reward using its own for with an independent replay buffer for each agent. Note that independent agents will still deviate due to the stochastic nature of the transition probabilities leading to the sampling of different actions .
In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions [38]. In this case, using for where the states and may be shared across agents in a cooperative manner. For simplicity, here we refer to the complete molecule and final state simply as the state. We investigated the following methods of cooperation (for additional detail and equations see Appendix B):
@@ 4.1 Molecule exploration benchmark
To address our proposed problem formulation to identify all hits in chemical space, we designed a new molecular exploration (MolExp) benchmark. Crucially, a ground truth is required to know that all solutions are identified 111Oracles that are more practically relevant, such as bioactivity prediction or docking, don’t contain a ground truth set of all possible high scoring solutions.. For this reason, we use the toy-task of multiple molecule rediscovery via optimization of a similarity scoring function. To ensure relevance to real-world discovery, each set of molecules are at minimum pre-clinical drug candidates bioactive against the same target. Moreover, to make the objective more challenging, the set of molecules are selected to be structurally dissimilar albeit possessing similar bioactivity. This phenomena is the converse of the similarity principle 222In drug discovery, the similarity principle states that similar molecules possess similar bioactivity . However, this does not prove the converse and many exceptions exist that are more difficult to model.. Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models [39, 40].
Each run is replicated on the benchmark 5 times with a default budget of 10,000 molecules per run, consistent with previous benchmarks [37].
@@
The benchmark consists of four tasks. Each task is based on a set of drug candidates and their shared biological target, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor drug candidates, beta-secretase 1 (BACE1) drug candidates, and epidermal growth factor drug candidates (EGFR). These tasks contain a set of 2, 3, 3, & 4 target molecules respectively, where the objective is to rediscover the full set of target molecules. To mimic a chemical space where distinct regions of reward signal exist, the oracle returns the similarity to the nearest target molecule as the reward . As CLMs act in string space, we chose a string-edit similarity function based on Levenshtein distance (MolExpL). However, we additionally report the results when using fingerprint similarity (MolExp) in the Appendix. A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix C.
@@
Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively . This metric focuses on the ability generate similar molecules to all objectives. Unless stated otherwise, we use reflecting the maximum similarity achieved. Molecular diversity is also measured by sphere exclusion diversity (SEDiv) at a sample size of 1,000 [41] which provides a more accurate representation of chemical space coverage than internal diversity [42, 30].
@@
We tested virtual screening, MolRL-MGPT [36], and multiple ACEGEN RL algorithms [23, 19] on their inherent ability for chemical exploration on the MolExpL benchmark (see Appendix D for implementation details). Note that MolRL-GPT was pre-trained on the same dataset for standardized comparison (see Appendix A for details). Table 1 shows that with a budget of 10,000 the ACEGENMolOpt algorithm achieved the highest score, therefore, we use this RL configuration for all further experiments. However, Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate), and that the best scores are due to improved efficiency in maximizing similarity to a single target molecule. Therefore, there is still considerable room for improvement. These results are also observed when using fingerprint similarity oracles (see subsection E.2).
@@
An estimated maximum performance can be calculated for ACEGENMolOpt RL configuration by training the agent to maximize only the similarity to each individual molecular target with the reward , (see Figure 14). This tests the agents ability to maximize similarity to each individual molecular target only, an expected prerequisite for maximizing similarity to multiple target molecules in the benchmark. Calculating the equivalent benchmark score results in an estimated maximum MolExpL score of 2.91.
@@
For reference, baseline algorithms were tested on the GuacaMol benchmark also with a fixed budget of 10,000 molecules. This shows ACEGENMolOpt achieving the highest score (see subsection E.3). However, a key difference is that 5/6 GuacaMol similarity and rediscovery tasks are already solved by ACEGENMolOpt with a score , whilst the MolExpL benchmark results in lower scores reflecting a higher difficulty of benchmark tasks.
@@ 4.2 Test-time training scaling for chemical exploration
We tested if scaling TTT with larger population sizes of independent RL agents would increase chemical exploration and improve benchmark performance. In this case, we chose the best-performing RL configuration-ACEGENMolOpt-and assigned a budget of 10,000 molecules to each agent in the population. 2(a) shows that MolExpL performance increases log-linearly from with 1 agent to with 128 agents, almost reaching a maximum score of 4. Log-linear scaling was also confirmed using the vanilla REINFORCE RL configuration, albeit with lower benchmark scores. Interestingly, when the population size reached 32 the estimated maximum score for this RL configuration was exceeded, indicating that the additional variance between independent agents overcomes the additionally reward signal ambiguity introduced from multiple molecular targets. Figure 17 shows that by increasing agents, there is more variance across the set of molecule targets. The AP task is an exception, where all 128 independent agents all maximize similarity to the same target molecule in the set. Similar results are also observed when using fingerprint similarity oracles subsection E.2. Additionally, similar scaling was achieved with up to 32 agents on the GuacaMol benchmark (subsection F.3), at which point performance on many tasks becomes saturated.
Increasing the number of agents, each with a budget of 10,000, increases the total budget. Therefore, we compared the alternative approach of scaling TTT via increasing the budget for a single RL agent. However, this did not result in similar MolExpL performance gain as shown in 2(b). Single RL agent performance saturates at a budget of 40,000 as the agent expectedly fails to continue exploring once one of the target molecules has been maximized (Figure 18). Moreover, 2(c) shows that molecular diversity decreases while scaling budget. Employing RND as a state-of-the-art exploration strategy, the most performant exploration strategy [35, 19], marginally improves performance and molecular diversity. Where Figure 19 shows that RND only results in fluctuations in similarity to one molecular target, as opposed to the desired behavior of switching from one molecular target in the set to another. We additionally tested a simple diversity filter (DF) that assigns a reward of 0 to non-unique canonical SMILES. However, the DF also failed to rescue TTT scaling in this dimension. Figure 3 exemplifies how 87 agents are required to solve the A2A task and recover all target molecules, while scaling one agent results in a focus only on the third target molecule.
@@ 4.3 Cooperative multi-agent RL for targeted chemical exploration
We investigated an extensive range of the different cooperative strategies on the MolExpL benchmark, including the MolRL-MGPT baseline with the same budget and number of agents, shown in 4(a). All of the strategies tested were similar or worse in performance, with the best being POPNORM, ENTS, and DIFFN. However, no strategy significantly outperformed independent agents as measured by a Bonferroni-corrected, one-tailed t-test. On the other hand, 4(b) shows that many cooperative strategies successfully resulted in corresponding increases in molecular diversity, namely RND, ENTS, CES, and DIFFS. This is usually anti-correlated with MolExpL score. This highlights the difficulty and complexity in achieving targeted exploration, where diversity appears to hinder learning, or is naive rather than targeted on on favorable regions of chemical space. These trends were even more prominent when using fingerprint similarity oracles subsection G.2.
@@ 4.4 Maximizing predicted A2A bioactivity
5(a) shows that scaling the number of independent RL agents increases the benchmark score, again more than scaling the budget of a single RL agent (5(b)). In other words, scaling RL agents increases the maximum similarity to the set of target molecules, and hence the chance of rediscovering the drug candidates. This represents increased chemical exploration of the reward landscape. This task highlights the translational benefit of TTT scaling with RL agents to real-world drug design objectives.
@@ 5 Conclusion
We did not investigate the effect of model architecture and training dataset in this work, which has been thoroughly investigated elsewhere. Instead we focus on maximizing chemical exploration efficiency given the same or similar starting policy. These results are expected to translate to any starting policy.
@@
The code used to generate this data is part of the ACEGEN repository https://github.com/Acellera/acegen-open. The code and training dataset for the MolExp benchmark is part of the MolScore repository https://github.com/MorganCThomas/MolScore which is integrated into ACEGEN.
@@ A.1 Datasets
MolExp benchmark The pre-training dataset for the MolExp(L) benchmark was curated from the ChEMBL34 database [44]. First, the database was filtered to ensure no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set . All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated. This resulted in a dataset of 1,711,022 unique molecules. Finally, the target molecules of the MolExp benchmark Appendix C were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
@@ A.2 ACEGEN
MolExp benchmark The CLM constituting the prior policy is a gated recurrent unit (GRU) network with an embedding of 256 and 3 layers of GRU cells with hidden dimension 512. This was trained with restricted SMILES randomization [45] for 10 epochs with a batch size of 128. This was implemented such that a random SMILES string augmentation was applied each time a dataset molecule was sampled during training. An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps. The learning rate, training loss, and SMILES validity during training is shown in Figure 6.

### v3

@@ 1 Introduction
Thorough exploration of chemical space is critical to maximize the potential of generative AI in drug design. Identifying structurally diverse molecules with similar bioactivity increases the likelihood of discovering backup series with distinct chemical and biological profiles, mitigating risks in drug development [4]. Moreover, exploring novel chemical regions can circumvent patented compounds [5]. However, many generative models, including CLMs, are susceptible to mode collapse when fine-tuned and depending on hyperparameters [6],particularly those reliant on machine learning models with limited applicability domains due to epistemic uncertainty [7, 8].
@@ 2.1 Problem definition
However, chemical space is vast, property predictors are inherently inaccurate with high uncertainty, and Pareto optimal solutions have different property profiles. Hence, multiple molecules exist across chemical space that are considered hits if the score is above a given threshold (or sometimes more simply the top- molecules).
We propose that the problem of generative molecular design should extend to finding all possible hits across chemical space , and postulate that this better aligns with practical requirements in drug discovery.
We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, even structurally unrelated molecular targets in chemical space (Figure 2.1, right). Previously published benchmarks don’t test this adequately, with objectives focused on a single narrow target, or a single broad target (Figure 2.1, left) and performance measured only by maximum achieved. Alternative evaluations have been proposed that also measure the number of diverse hits [30, 31]. Although they are an improvement in benchmark evaluation, there are still some limitations to this approach. Firstly, rewarding diverse hits are not suitable for many tasks converging to single solutions, hence why Renz et al. 2024 focus only on 3/23 tasks from MolOpt. Secondly, tasks where diverse hits are suitable, such as property prediction where multiple modes exist, are typically trivial. In some cases, maximum performance is achieved and hence no statistical differences can be observed [37]. Lastly, when measuring the number of diverse hits it is not known how many possible diverse hits exist in solution space i.e., there is no ground truth. We propose to add a ground truth via a limited number of known modes in solution space. Moreover, ensuring the modes are distinct and structurally unrelated in a chemical sense, increases the difficulty of the benchmark and inherent exploration required, enabling the measurement of statistically significant improvements.
@@ 2.2 Reinforcement learning with chemical language models
CLMs utilize a language representation, commonly SMILES [9], to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models. For example, SMILES represent the molecular structure of paracetamol with the string “CC(=O)Nc1ccc(O)cc1". These models operate autoregressively, predicting the next token in a sequence based on the context provided by previous tokens. The process can be formalized as maximizing the conditional probability of each token given the preceding sequence , where denotes the tokens up to . The model is trained using the cross-entropy loss function which is equivalent to the negative log-likelihood Equation 2.4. Where is the length of the -th sequence, and are the model parameters.
@@
We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the trajectory probability by prior policy denoted as with a coefficient and by adding an exponent , as proposed by Thomas et al. [20]. These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape. In addition to reshaping the reward, we implement experience replay to augment on-policy data with off-policy data at each iteration. A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward. This additionally improves learning efficiency [38, 20].
@@ 2.3 Cooperative reinforcement learning
In the independent case, each agent still aims to maximize its cumulative reward using its own for with an independent replay buffer for each agent. Note that independent agents will still deviate due to the stochastic nature of policy sampling from the transition probabilities leading to the sampling of different actions .
In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions [39]. In this case, using for where the states and may be shared across agents in a cooperative manner. We investigated the following methods of cooperation (for additional detail and equations see Appendix F):
@@ 3.1 Molecule exploration benchmark
To address our proposed problem formulation to identify all hits in chemical space, we designed a new molecular exploration (MolExp) benchmark. Crucially, a ground truth is required to know that all solutions are identified 111Oracles that are more practically relevant, such as bioactivity prediction or docking, don’t contain a ground truth set of all possible high scoring solutions.. For this reason, we use the toy-task of multiple molecule rediscovery via optimization of a similarity scoring function. To ensure relevance to real-world discovery, each set of molecules are at minimum pre-clinical drug candidates bioactive against the same target. Moreover, to make the objective more challenging, the set of molecules are selected to be structurally dissimilar albeit possessing similar bioactivity. This phenomena is the converse of the similarity principle 222In drug discovery, the similarity principle states that similar molecules possess similar bioactivity . However, this does not prove the converse and many exceptions exist that are more difficult to model.. Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models [40, 41]. To test translation of the benchmark to more practically applicable objectives, we also test chemical exploration when optimizing the predicted A2A bioactvity which we call MolExpBio.
Each run is replicated on the benchmark 5 times.
@@
The benchmark consists of four tasks. Each task is based on a set of drug candidates and their shared biological target, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor drug candidates, beta-secretase 1 (BACE1) drug candidates, and epidermal growth factor drug candidates (EGFR). These tasks contain a set of 2, 3, 3, & 4 target molecules respectively, where the objective is to rediscover the full set of target molecules. This is similar in prospective drug discovery projects where 1 drug candidate will be progressed, but ideally with a small number of backup candidates. To mimic a chemical space where distinct regions of reward signal exist, the oracle returns the similarity to the nearest target molecule as the reward . As CLMs act in string space, we chose a string-edit similarity function based on Levenshtein distance (MolExpL). However, we additionally report the results when using fingerprint similarity (MolExp) in the Appendix. A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix B. For the MolExpBio experiment, the oracle is the predicted probability of A2A bioactivity, predicted by a Random Forest classification model trained on ECFP4 fingerprints of actives and inactives extracted from ChEMBL [42] made available in MolScore [43].
@@
Performance is measured by calculating the product of the maximum similarity achieved to each target molecule . This metric focuses on the ability to generate similar molecules to all objectives. Molecular diversity is also measured by sphere exclusion diversity (SEDiv) at a sample size of 1,000 [44] which provides a more accurate representation of chemical space coverage than internal diversity [45, 31]. Note this is equivalent to the later published #Circles [45] metric at a Tanimoto threshold of 0.65 and normalized by the sample size.
@@
We tested virtual screening, MolRL-MGPT [37], and multiple ACEGEN RL algorithms [24, 20] on their inherent ability for chemical exploration on the MolExpL benchmark (see Appendix C for implementation details). Note that MolRL-GPT was pre-trained on the same dataset for standardized comparison (see Appendix A for details). Table 3.1 shows that with a budget of 10,000 the ACEGENMolOpt algorithm achieved the highest score, therefore, we use this RL configuration for all further experiments. However, Figure D.1 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate), and that the best scores are due to improved efficiency in maximizing similarity to a single target molecule. Therefore, there is still considerable room for improvement. These results are also observed when using fingerprint similarity oracles (see subsection D.2).
@@
An estimated maximum performance can be calculated for ACEGENMolOpt RL configuration by training the agent to maximize only the similarity to each individual molecular target with the reward , (see Figure D.2). This tests the agents ability to maximize similarity to each individual molecular target only, an expected prerequisite for maximizing similarity to multiple target molecules in the benchmark. Calculating the equivalent benchmark score results in an estimated maximum MolExpL score of 2.91.
@@
For reference, baseline algorithms were tested on the GuacaMol benchmark also with a fixed budget of 10,000 molecules. This shows ACEGENMolOpt achieving the highest score (see subsection D.3). However, a key difference is that 5/6 GuacaMol similarity and rediscovery tasks are already solved by ACEGENMolOpt with a score , whilst the MolExpL benchmark results in lower scores reflecting a higher difficulty of benchmark tasks.
@@ 3.2 Test-time training scaling for chemical exploration
We tested if scaling TTT with larger population sizes of independent RL agents would increase chemical exploration and improve benchmark performance. In this case, we chose the best-performing RL configuration-ACEGENMolOpt-and assigned a budget of 10,000 molecules to each agent in the population. 1(a) shows that MolExpL performance increases log-linearly from with 1 agent to with 128 agents, almost reaching a maximum score of 4. Log-linear scaling was also confirmed using the vanilla REINFORCE RL configuration, albeit with lower benchmark scores. Interestingly, when the population size reached 32 the estimated maximum score for this RL configuration was exceeded, indicating that the additional variance between independent agents overcomes the additionally reward signal ambiguity introduced from multiple molecular targets. Figure E.1 shows that by increasing agents, there is more variance across the set of molecule targets. The AP task is an exception, where all 128 independent agents all maximize similarity to the same target molecule in the set. Similar results are also observed when using fingerprint similarity oracles subsection D.2. Additionally, similar scaling was achieved with up to 32 agents on the GuacaMol benchmark (subsection E.3), at which point performance on many tasks becomes saturated.
Increasing the number of agents, each with a budget of 10,000, increases the total budget. Therefore, we compared the alternative approach of scaling TTT via increasing the budget for a single RL agent. However, this did not result in similar MolExpL performance gain as shown in 1(b). Single RL agent performance saturates at a budget of 40,000 as the agent expectedly fails to continue exploring once one of the target molecules has been maximized (Figure E.2). Moreover, 1(c) shows that molecular diversity decreases while scaling budget. Employing RND as a state-of-the-art exploration strategy, the most performant exploration strategy [36, 20], marginally improves performance and molecular diversity. Where Figure E.3 shows that RND only results in fluctuations in similarity to one molecular target, as opposed to the desired behavior of switching from one molecular target in the set to another. We additionally tested a simple diversity filter (DF) that assigns a reward of 0 to non-unique canonical SMILES. However, the DF also failed to rescue TTT scaling in this dimension. Figure 3.2 exemplifies how 87 agents are required to solve the A2A task and recover all target molecules, while scaling one agent results in a focus only on the third target molecule.
@@ 3.3 Cooperative multi-agent RL for targeted chemical exploration
We investigated an extensive range of the different cooperative strategies on the MolExpL benchmark, including the MolRL-MGPT baseline with the same budget and number of agents, shown in 3(a). All of the strategies tested were similar or worse in performance, with the best being POPNORM, ENTS, and DIFFN. However, no strategy significantly outperformed independent agents as measured by a Bonferroni-corrected, one-tailed t-test. On the other hand, 3(b) shows that many cooperative strategies successfully resulted in corresponding increases in molecular diversity, namely RND, ENTS, CES, and DIFFS. This is usually anti-correlated with MolExpL score. This highlights the difficulty and complexity in achieving targeted exploration, where additional diversity appears to hinder learning, or is naive rather than targeted on favorable regions of chemical space. These trends were even more prominent when using fingerprint similarity oracles subsection G.2.
@@ 3.4 Maximizing predicted A2A bioactivity
4(a) shows that scaling the number of independent RL agents increases the benchmark score, again more than scaling the budget of a single RL agent (4(b)). In other words, scaling RL agents increases the maximum similarity to the set of target molecules, and hence the chance of rediscovering the drug candidates. This represents increased chemical exploration of the reward landscape. This task highlights the translational benefit of TTT scaling with RL agents to real-world drug design objectives.
@@ 4 Conclusion
We did not investigate the effect of model architecture and training dataset in this work, which has been thoroughly investigated elsewhere [12, 46]. Instead we focus on maximizing chemical exploration efficiency given the same or similar starting policy. These results are expected to translate to any starting policy.
@@
All code is openly available under an MIT license. ACEGEN is available on GitHub at https://github.com/Acellera/acegen-open, whilst the benchmark is implemented in MolScore available on GitHub at https://github.com/MorganCThomas/MolScore. Note the MolExp pretraining dataset is also available in the MolScore repository.
@@ A.1 Datasets
MolExp benchmark The pre-training dataset for the MolExp(L) benchmark was curated from the ChEMBL34 database [47]. First, the database was filtered to ensure no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set . All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated. This resulted in a dataset of 1,711,022 unique molecules. Finally, the target molecules of the MolExp benchmark Appendix B were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
@@ A.2 ACEGEN
MolExp benchmark The CLM constituting the prior policy is a gated recurrent unit (GRU) network with an embedding of 256 and 3 layers of GRU cells with hidden dimension 512. This was trained with restricted SMILES randomization [48] for 10 epochs with a batch size of 128. This was implemented such that a random SMILES string augmentation was applied each time a dataset molecule was sampled during training. An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps. The learning rate, training loss, and SMILES validity during training is shown in Figure A.1.